OpenStudy (anonymous):

find the solution of equation3x+2y=4 and 5x-4y=3

OpenStudy (shubhamsrg):

Any attempt ?

OpenStudy (anonymous):

3x+2y=4---------(1 5x-4y=3---------(2 Multiplying eq. (1 by 2 and then adding it with eq(2, we find 6x+4y=8---------(1 5x-4y=3---------(2 ---------------------- 11x = 11 hence x= 11/11 i.e. x= 1 substituting x=1 in eq(1 we find 3*1+2y=4 3+2y =4 2y =4-3 2y =1 i.e. y= 1/2 =0.5 Hence (x, y) =( 1, 1/2) are the required solutions of the given system of equations @marey65

OpenStudy (anonymous):

thanks I have the same answer you are correct

OpenStudy (anonymous):

5x - 4y = 3 5x = 4y + 3 x = (4/5)y + (3/5) 3x = (12/5)y + (9/5) Now that we know what 3x is in terms of y, we can put it into the equation 3x + 2y = 4. Then simplify. (12/5)y + (9/5) + 2y = 4 2.4y + 1.8 + 2y = 4 4.4y + 1.8 = 4 4.4y = 2.2 y = .5 (1/2) Now put the y value into the equations to get x. 3x + 2(.5) = 4 3x + 1 = 4 3x = 3 x = 1 5x - 4(.5) = 3 5x - 2 = 3 5x = 5 x = 1 So your answer is: x = 1, y = .5